A Carnot cycle heat engine operates between 172 K and 335 K. B Its efficiency is?
1
Expert's answer
2019-01-21T11:10:56-0500

We can find the efficiency of the Carnot heat engine from the formula:

"\eta = (1 - \frac{T_C}{T_H})\cdot 100\%,"

here,

"T_C"

is the absolute temperature of the cold reservoir,

"T_H"

is the absolute temperature of the hot reservoir.

Then, we get:

"\eta = (1 - \frac{T_C}{T_H})\cdot 100\% = (1 - \frac{172 K}{335 K})\cdot 100\% = 48.6\%."

Answer:

"\eta = 48.6\%"

.